In a triangle with angles measuring a, b, and c degrees, the mean of b and c is a. What is the value of a?
Answer by ReadingBoosters(3246)

Knowing that a+b+c=180
and (b+c)/2=a
Then,
(180-a)/2 = a
Multiply both sides by 2
180 - a = 2a
Group and combine like terms.
180 = 3a
Divide both sides by 3
a = 60 degrees
...................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.